OpenStudy (anonymous):

what is equivalent to 4/9

OpenStudy (mathmale):

In other words, what's the decimal fraction equivalent to 4/9? How many decimal place accuracy do you want?

OpenStudy (anonymous):

Hi ! \[\frac{ 4 }{ 9 }=\left(\begin{matrix}2 \\ 3\end{matrix}\right)^{2}\]

OpenStudy (texaschic101):

if you multiply the top and bottom by the same number, you will end up with an equivalent fraction. Example : 4/9 x 2/2 = 8/18 (this is equivalent) 4/9 x 3/3 = 12/27 (this is equivalent) 4/9 x 4/4 = 16/36 (this is equivalent) I think you get my point

OpenStudy (anonymous):

\((\frac23)(\frac23)=\frac49\)

OpenStudy (texaschic101):

it could be lots of things....we don't even have answer choices

OpenStudy (mathmale):

And therefore (2/3)^2 is also equivalent to 4/9.

OpenStudy (mathmale):

Better find out, in one way or another, precisely what your teacher expects you to be able to do in this kind of math problem solving.

OpenStudy (anonymous):

8/18